OpenStudy (anonymous):

help this simply derivative question: f(x)= sqrt(5x)

OpenStudy (anonymous):

i rewrote it: (5x)^1/2 then multiplied and take away one from the exponent: (5x/2) ^-1/2 then i move the negative exponent in the denominator: 5x/2 sqrt x)

OpenStudy (anonymous):

how come 5 is also a square root?

OpenStudy (anonymous):

what step am i missing?

OpenStudy (anonymous):

\[ \sqrt{5x} = \sqrt{5}\sqrt{x} \]

OpenStudy (anonymous):

So the derivative is \[ \sqrt{5}\frac{1}{2\sqrt{x}} \]

OpenStudy (anonymous):

@mathcalculus Make sense?

OpenStudy (anonymous):

\[\eqalign{ &f(x)=\sqrt{5x} \\ &\\ &\\ &\frac{d}{dx}f(x)=\frac{d}{dx}\sqrt{5x}=\frac{d}{dx}\sqrt{5}\sqrt{x}=\sqrt{5}\frac{d}{dx}\sqrt{x}=\sqrt{5}\frac{1}{2\sqrt{x}}=\frac{\sqrt{5}}{2\sqrt{x}}=\frac{\sqrt{5}}{\sqrt{4x}}=\sqrt{\frac{5}{4x}} \\ }\]

OpenStudy (anonymous):

OOPS lol got cut off. wait...

OpenStudy (anonymous):

\[\eqalign{ &f(x)=\sqrt{5x} \\ &\\ &\\ &\eqalign{\frac{d}{dx}f(x) &=\frac{d}{dx}\sqrt{5x} \\ &=\frac{d}{dx}\sqrt{5}\sqrt{x} \\ &=\sqrt{5}\frac{d}{dx}\sqrt{x} \\ &=\sqrt{5}\frac{1}{2\sqrt{x}} \\ &=\frac{\sqrt{5}}{2\sqrt{x}} \\ &=\frac{\sqrt{5}}{\sqrt{4x}} \\ &=\sqrt{\frac{5}{4x}} \\} \\ }\]

OpenStudy (anonymous):

@wio so you can't re-write 5 as 5^(1/2)?

OpenStudy (anonymous):

Well, you can! But it would still be a constant.

OpenStudy (anonymous):

?